OpenStudy (anonymous):

Rotation of a Hyperbola Rotate the axes to eliminate the xy term and write the equation xy + 2 = 0 in standard form and sketch its graph.

OpenStudy (experimentx):

let X = xcos(A) + y Sin(A) Y = ycos(A) - xSin(A) find the value of A such that xy term vanishes.

OpenStudy (anonymous):

what is A?

OpenStudy (experimentx):

A is an angle of rotation.

OpenStudy (experimentx):

put those values in x and y, multiply them, can separate XY term, and equate the coefficients of xy ... and get the value of A

OpenStudy (anonymous):

so what would i put into x and y? and what does the X and Y mean are thet different from the lower case x and y?

OpenStudy (experimentx):

yes big X and big Y

OpenStudy (anonymous):

hold up what do i put into x and y?

OpenStudy (experimentx):

(Xcos(A) + Y Sin(A))(Ycos(A) - XSin(A)) + 2 = 0

OpenStudy (anonymous):

im so sorry for asking dumb questions but i still dont get what i put into the X and Y...

OpenStudy (experimentx):

just multiply the above expression and simplify.

OpenStudy (experimentx):

you have already put ... not you but me. x*y+2 = 0 (Xcos(A) + Y Sin(A))*y + 2 = 0; (Xcos(A) + Y Sin(A))(Ycos(A) - XSin(A)) + 2 = 0;

OpenStudy (anonymous):

how do i solve for X, Y or A?

OpenStudy (experimentx):

you don't need to solve your X,Y Just solve for A, equating coefficinets of XY to zero.

OpenStudy (anonymous):

so i simplified it and got XY(cos(A)) - X^2(cos(A))(sin(A)) + y^2(sin(A))(cos(A)) - XY(sin(A))^2= -2 what do i do now?

OpenStudy (experimentx):

coeffiecitnt of XY must be zero ... from there get the value of A

OpenStudy (anonymous):

-X^2(cos(A))(sin(A)) + y^2(sin(A))(cos(A))= -2 so would it look like this then? if it does then i dont know where to go from here.....sorry for bothering you so much.. :)

OpenStudy (experimentx):

no get A by equationg the coefficein of XY = zero

OpenStudy (anonymous):

how would you do that?

OpenStudy (experimentx):

where did you put your xy terms??

OpenStudy (anonymous):

i thought they were zero so i took then off cuase that what i thought you ment...

OpenStudy (anonymous):

okay so starting over did i do this equation right? XY(cos(A)) - X^2(cos(A))(sin(A)) + y^2(sin(A))(cos(A)) - XY(sin(A))^2= -2

OpenStudy (experimentx):

XY(cos(A)) - XY(sin(A))^2 <--- it's coefficient must be zero.

OpenStudy (anonymous):

aight THANKS SO MUCH i think the standard form would be (y/4)-(x/4)=1and its tilted at a 45 degree angle....

OpenStudy (experimentx):

well, that is a line, not a hyperbola ... and your original curve was rectangular huperbola.

OpenStudy (experimentx):

anyway welcome.

OpenStudy (anonymous):

no thts standard form of a hyperbola....its definity not a line